OpenStudy (anonymous):

find the domain of the composite function f o g f(x)=3/x-10 g(x)=sqrt(x-3)

OpenStudy (anonymous):

first off you need the domain of \(g\) you got that?

OpenStudy (anonymous):

if not let me know i am just asking

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then you have \[f\circ g(x)=\frac{3}{\sqrt{x-3}-10}\]

OpenStudy (anonymous):

you already know that \(x\geq 3\) for sure now you have to make sure that you are not dividing by zero, solve \[\sqrt{x-3}-10=0\] for \(x\) to get another restriction

OpenStudy (anonymous):

x-5, x+2

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

\[\sqrt{x-3}=10\\ x-3=100\\ x=103\] is what i get

OpenStudy (anonymous):

how 10 become 100

OpenStudy (anonymous):

i squared both sides